A spur gear having module 7 mm. The smallest teeth numbers are 10. If the speed is 

1300 r.p.m. Then find out the pitch line velocity


1
Expert's answer
2021-02-16T07:29:53-0500

pitch line velocity = "Vp = \frac{\pi D1 N1}{60}"

"Vp = \frac{\pi\times 43.2\times 1300}{60\times 10}"


= 293